# The Race between Technology and Woman:

### v8

@@ 3.4 Effects of technological change
This equation is a multi-sector, multi-factor general equilibrium version of the Krusell, Ohanian, Rios-Rull and Violante (2000) model. It implies that the direction and magnitude of the effect of capital-embodied technological progress on the skill wage gap depend on the presence and degree of capital–skill complementarity. The skill wage gap can increase with capital-embodied technological progress if capital equipment is more complementary to skilled labor than to unskilled labor (i.e., $\epsilon_{\ell_{u}k_{e}}^{\mathcal{F}}>\epsilon_{\ell_{h}k_{e}}^{\mathcal{F}}$). The rate of the increase becomes greater as the difference between the two elasticities or the income share of capital equipment increases. The effect of sector-specific disembodied technological change on the skill wage gap can be divided into two parts. The first part in the square bracket takes a similar form to the effect of capital-embodied technological change and captures the effect through a change in capital equipment. However, this has the opposite sign to that of the first term because sector-specific disembodied technological progress reduces the need for the use of capital equipment. The second part is related to the direct effect denoted by $\psi^{\ell}$ in equation (19). The sign of the first part is negative in the presence of capital–skill complementarity, whereas that of the second part is likely to be positive. In this setting, sector-specific disembodied technological progress is unlikely to account for the widening of the skill wage gap.
3.5 Extension
3.5.1 Endogenous labor supply
Proposition 5.
$d\ln\left(\frac{w_{f}}{w_{g}}\right)=\sum_{h=1}^{F_{k}}\left(\Xi_{\left(f,h\right)}^{q}-\Xi_{\left(g,h\right)}^{q}\right)d\ln q_{h}+\sum_{n=1}^{N}\left(\Xi_{\left(f,n\right)}^{A}-\Xi_{\left(g,n\right)}^{A}\right)d\ln A_{n}\quad\text{for }f,g=1,\ldots,F_{\ell},$ (23)
where $\Xi_{\left(f,g\right)}^{q}$ is the $fg$-th element of the $F_{\ell}\times F_{k}$ matrix $\Xi^{q}$, and $\Xi_{\left(f,n\right)}^{A}$ is the $fn$-th element of the $F_{\ell}\times N$ matrix $\Xi^{A}$. The matrices $\Xi^{q}$ and $\Xi^{A}$ are described in Appendix A.1. Changes in labor shares due to technological change are given by
$d\ln\Lambda_{\ell_{f}}=\sum_{h=1}^{F_{k}}\left\{\sum_{g\neq f}^{F}\left(1-\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}\Xi_{\left(f,h\right)}^{q}+\sum_{g\neq f}^{F_{\ell}}\left[-\left(1-\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}+\sum_{s\neq f,g}^{F}\left(\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}-\epsilon_{\ell_{s}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{s}}\right]\Xi_{\left(g,h\right)}^{q}\right.\\ \left.-\left[-\left(1-\epsilon_{\ell_{f}k_{h}}^{\mathcal{C}}\right)\Lambda_{k_{h}}+\sum_{s\neq f,h}^{F}\left(\epsilon_{\ell_{f}k_{h}}^{\mathcal{C}}-\epsilon_{\ell_{s}k_{h}}^{\mathcal{C}}\right)\Lambda_{\ell_{s}}\right]\right\}d\ln q_{h}+\sum_{n=1}^{N}\left\{\Phi_{\left(f,n\right)}^{w}+\sum_{g\neq f}^{F}\left(1-\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}\Xi_{\left(f,n\right)}^{A}\right.\\ \left.+\sum_{g\neq f}^{F_{\ell}}\left[-\left(1-\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}+\sum_{s\neq f,g}^{F}\left(\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}-\epsilon_{\ell_{s}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{s}}\right]\Xi_{\left(g,n\right)}^{A}\right\}d\ln A_{n}\quad\text{for }f=1,\ldots,F_{\ell}.$ (24)
Equations (23) and (24) take similar forms to equations (21) and (22), respectively, and nest them as special cases with zero elasticity of labor supply. Most elements of the matrix $\Xi^{q}$ consist of the aggregate elasticities of substitution $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}$ only or both the aggregate elasticities of substitution and the aggregate elasticity of labor supply $1/\gamma$, and the others consist of the factor shares of income $\Lambda_{\ell_{f}}$, as detailed in the appendix. Equations (23) and (24) show that the aggregate elasticities of substitution and the aggregate elasticity of labor supply are sufficient statistics to quantify the effects of technological change on the relative wages and income shares of different types of labor if the income shares of each factor are known.
Corollary 8.
3.5.2 Imperfect labor mobility
Proposition 6.
$d\ln\left(\frac{\overline{w}_{f}}{\overline{w}_{g}}\right)=\sum_{h=1}^{F_{k}}\sum_{n=1}^{N}\left(\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\Omega_{f\left(n,h\right)}^{q}-\frac{\Lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\Omega_{g\left(n,h\right)}^{q}\right)d\ln q_{h}\\ +\sum_{n=1}^{N}\sum_{m=1}^{N}\left(\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\Omega_{f\left(n,m\right)}^{A}-\frac{\Lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\Omega_{g\left(n,m\right)}^{A}\right)d\ln A_{m}\quad\text{for }f,g=1,\ldots,F_{\ell},$ (25)
where $\Omega_{f(n,g)}^{q}$ is the $ng$-th element of the $N\times F_{k}$ matrix $\Omega_{f}^{q}$, and $\Omega_{f(n,m)}^{A}$ is the $nm$-th element of the $N\times N$ matrix $\Omega_{f}^{A}$. The matrices $\Omega_{f}^{q}$ and $\Omega_{f}^{A}$ are described in Appendix A.1. Changes in labor shares due to technological change are given by
$d\ln\Lambda_{\ell_{f}}=\sum_{h=1}^{F_{k}}\sum_{n=1}^{N}\sum_{m=1}^{N}\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\left(\sum_{g=1}^{F_{\ell}}\Pi_{UL,fg\left(n,m\right)}^{w}\Omega_{g\left(m,h\right)}^{q}-\Pi_{UR,fh\left(n,m\right)}^{w}\right)d\ln q_{h}\\ +\sum_{n=1}^{N}\sum_{m=1}^{N}\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\left(\sum_{g=1}^{F_{\ell}}\sum_{l=1}^{N}\Pi_{UL,fg\left(n,l\right)}^{w}\Omega_{g\left(l,m\right)}^{A}+\Theta_{f\left(n,m\right)}^{w}\right)d\ln A_{m}\quad\text{for }f=1,\ldots,F_{\ell},$ (26)
where $\Pi_{UR,fg(n,m)}^{w}$ is the $nm$-th element of the $N\times N$ matrix $\Pi_{UR,fg}^{w}$, $\Pi_{UL,fg(n,m)}^{w}$ is the $nm$-th element of the $N\times N$ matrix $\Pi_{UL,fg}^{w}$, and $\Theta_{f}^{w}$ is the $N\times N$ matrix whose $nm$-th element is given by $\Theta_{f(n,m)}^{w}=-[1-1/(1-\eta)](I_{(n,m)}-\zeta_{m})$. The matrices $\Pi_{UR,fg}^{w}$ and $\Pi_{UL,fg}^{w}$ are described in Appendix A.1.
Most elements of the matrix $\Omega_{f}^{q}$ consist of the sectoral elasticities of substitution defined as $\epsilon_{\ell_{fn}\ell_{gm}}^{\mathcal{C}}=\partial\ln(\ell_{fn}/\ell_{gm})/\partial\ln(w_{gm}/w_{fn})$, and the others are the factor shares of income $\Lambda_{\ell_{f}n}$, as detailed in the appendix. All elements of the matrices $\Pi_{UR,fg}^{w}$ and $\Pi_{UL,fg}^{w}$ consist of the sectoral elasticities of substitution $\epsilon_{\ell_{fn}\ell_{gm}}^{\mathcal{C}}$ and the factor shares of income $\Lambda_{\ell_{f}n}$. Equations (25) and (26) show that the sectoral elasticities of substitution are sufficient statistics to quantify the effects of capital-embodied technological change on the relative wages and income shares of different types of labor if the income shares of each factor are known.
@@ 4 Data
We start this section by describing the data sources, sample, and variables used in the analysis. We then document trends in the relative prices, relative quantities, and income shares of factors. We end this section by presenting preliminary estimates of substitutability between ICT capital and the four types of labor.
@@ 4.1 Data sources, sample, and variables
The data used in the analysis are mainly from the EU KLEMS database, which collects detailed and internationally comparable information on the prices and quantities of capital and labor in major OECD countries. Our analysis is based on the March 2008 version because it contains the longest time series covering the 1980s, during which there were dramatic changes in technology and inequality. For the years 1980 to 2005, we include all countries and years in the sample for which required data are available. Our sample comprises 11 OECD countries: Australia, Austria, the Czech Republic, Denmark, Finland, Germany, Italy, Japan, the Netherlands, the United Kingdom, and the United States. The economy of each country is divided into five sectors: primary, manufacturing, utilities, construction, and services. [^4] The services sector includes nine categories: wholesale and retail trade; hotels and restaurants; transport and storage, and communication; financial intermediation; real estate, renting, and business activities; public administration and defense and compulsory social security; education; health and social work; and other community and social and personal services. The sample comprises four non-primary sectors and includes 258 country-year observations for each sector.
Labor is divided into skilled and unskilled labor, each of which is further divided into male and female labor. Skilled labor consists of workers who have completed college, and unskilled labor consists of workers who do not enter or complete college. We calculate wages by gender and skill by dividing the total labor compensation by the total hours worked. When we calculate wages and hours worked, we adjust for changes in the age and education composition of the labor force over time. Appendix A.3.1 provides the details of the adjustment procedure.
Capital is divided into ICT and non-ICT capital. ICT capital consists of computing equipment, communications equipment, and software. Non-ICT capital consists of transport equipment, other machinery and equipment, non-residential structures and infrastructures, residential structures, and other assets. We calculate the rental price of capital (Jorgenson, 1963), based on either the internal or external rate of return. The former approach has the advantage of maintaining consistency between national income and production accounts, whereas the latter approach has the advantage of not requiring the assumption of competitive markets. We employ the latter approach to examine the robustness of our results. Appendix A.3.2 provides the details of the calculation procedure.
@@ 4.2 Factor prices, quantities, and shares
During the period between the years 1980 and 2005, there was a difference in the trends of the gender wage gap between skilled and unskilled workers in OECD countries (Figure 1a). The male–female wage gap declined among unskilled workers, but not among skilled workers. At the same time, there was a difference in the trends of the skill wage gap between male and female workers in OECD countries (Figure 1b). The skilled–unskilled wage gap increased among male workers after a slight drop in the early 1980s, while it did not increase among female workers except for the period between the years 1987 and 1995. Taken together, the rate of decrease in the male–female wage gap was greater among unskilled workers than among skilled workers, while the rate of increase in the skilled–unskilled wage gap was greater among male workers than among female workers. This finding is perhaps surprising but not inconsistent with the fact that the rate of decline in the gender wage gap was greater at the middle or bottom than at the top of the wage distribution in the United States from the 1980s to the 2000s (Blau and Kahn, 2017). The patterns of changes in gender and skill premia are more visible in the United States (Figures 2a and 2b).
@@
Turning to the relative quantities of labor, there was a decline in the quantity of male labor relative to female labor, and an increase in the quantity of skilled labor relative to unskilled labor. The rate of decline in the relative quantity of male labor was much greater for skilled labor than for unskilled labor (Figure 3a). At the same time, the rate of increase in the relative quantity of skilled labor was much greater for female labor than for male labor (Figure 3b). These observations suggest that the relative supply of labor alone could not explain the differences in the trends of the gender wage gap and the skill wage gap.
@@
Figure 4: ICT and non-ICT capital
(a) Rental Prices
@@
Notes: The rental prices of ICT and non-ICT capital are denoted by $r_{i}$ and $r_{o}$, respectively, while the quantities of ICT and non-ICT capital are denoted by $k_{i}$ and $k_{o}$, respectively. All the series are logarithmically transformed and normalized to zero in the year 1980. The 1980 values of $r_{i}$ and $r_{o}$ are 1.12 and 0.09, respectively. The 1980 values of $k_{i}$ and $k_{o}$ are 31.14 and 2388.77, respectively.
The trends in the rental prices of capital differ significantly between ICT and non-ICT capital (Figure 4a). The rental price of ICT capital fell dramatically, but that of non-ICT capital remained almost unchanged. Meanwhile, the quantities of both ICT and non-ICT capital increased. However, the rate of increase in ICT capital was far greater than that in non-ICT capital (Figure 4b). These observations indicate substantial progress in ICT during the period. Appendix A.3.3 shows that even if non-ICT equipment is distinguished from non-ICT structures, there is no significant difference in the trends of the prices and quantities between them.
The direction and magnitude of changes in labor shares vary by gender and skill (Figure 5). The income shares of male and female unskilled labor decreased, while the income shares of male and female skilled labor increased. The magnitude of the decline in the income share of male unskilled labor is significantly greater than that of the changes in the income shares of the other three types of labor. These observations indicate that the decline in the labor share is attributable to the fall in the income share of male unskilled labor.
@@ 4.3 Factor substitution
Before we specify the sectoral production function, we conduct a preliminary analysis to acquire a general idea of substitutability between ICT capital and the four types of labor. For labor of type $f$ in sector $n$, we consider the following equation:
@@
where $\alpha_{nj}$ is a country-specific unobserved effect, $\nu_{njt}$ is the error term, and $j$ and $t$ are indices for countries and years. The parameter of interest, $\varepsilon_{f,n}$, captures the percentage change in the capital–labor ratio resulting from a one percent change in the relative price of labor to capital, and thus, represents the degree of substitution between ICT capital and labor of type $f$ in sector $n$.
When we estimate equation (27), we deal with two econometric issues: (i) unobserved heterogeneity and (ii) simultaneity. First, we control for country-specific unobserved effects by first differencing. In doing so, we take 5-year differences to address the measurement error problem. Then, we use instrumental variables to solve the simultaneity problem between prices and quantities. Because we are interested in the extent to which the relative demand for labor changes in response to a change in the relative price of labor, we use two supply shifters as instrumental variables. The first instrument is the lagged population size of each demographic group. The second instrument is the lagged price of ICT investment relative to consumption, which measures the level of investment-specific technology, as shown in equation (7).
Table 4.3 presents the ordinary least squares (OLS) and first-difference instrumental variable (IV) estimates of $\varepsilon_{f,n}$ for female skilled labor ($f=fh$), male skilled labor ($f=mh$), female unskilled labor ($f=fu$), and male unskilled labor ($f=mu$) in manufacturing, utilities, construction, and services sectors. Both estimates indicate that the degrees of substitution between ICT capital and the four types of labor are smaller in the order of female skilled, male skilled, female unskilled, and male unskilled labor for all sectors (i.e., $\varepsilon_{fh,n}<\varepsilon_{mh,n}<\varepsilon_{fu,n}<\varepsilon_{mu,n}$ for all $n$). The first-difference IV estimates differ by the type of labor more than the OLS estimates. In both cases, the differences in the estimates by the type of labor are more significant for manufacturing, utilities, and services sectors than for the construction sector. Similar results are obtained for the 1980–1995 and 1990–2005 periods. The degrees of substitution between ICT capital and unskilled labor become smaller from the former to the latter period, but ICT capital is more substitutable with male and female unskilled labor than with male and female skilled labor. The results described thus far are reinforced by the Anderson and Rubin (1949) confidence sets robust to weak instruments. The first-difference IV estimates remain largely unchanged regardless of whether the lagged population size or the lagged investment price is used as an instrument. Overall, ICT capital is more complementary not only to skilled labor than unskilled labor but also to female labor than male labor.
{sidewaystable}
Substitutability between ICT capital and the four types of labor
OLS First-difference IV1 First-difference IV2
@@
Manufacturing 0.737 1.092 1.135 1.225 0.580 1.189 1.540 1.618 0.562 0.986 1.305 1.339
(0.136) (0.144) (0.179) (0.154) {0.392, 0.878} {1.054, 1.400} {1.301, 1.835} {1.392, 1.913} {0.373, 0.751} {0.832, 1.160} {1.050, 1.607} {1.113, 1.604}
Utilities 0.369 0.869 0.941 1.207 0.487 1.118 1.298 1.541 0.482 0.975 1.156 1.319
(0.234) (0.159) (0.058) (0.143) {0.352, 0.702} {0.883, 1.551} {1.065, 1.689} {1.269, 2.040} {0.347, 0.647} {0.696, 1.339} {0.886, 1.504} {1.029, 1.692}
Construction 0.728 1.016 1.015 1.004 1.072 1.809 1.815 2.032 0.923 1.373 1.456 1.583
(0.338) (0.445) (0.201) (0.254) {0.321, 2.408} {1.079, 3.081} {1.122, 2.835} {1.340, 3.132} {0.454, 1.558} {0.874, 2.053} {0.970, 2.123} {1.092, 2.277}
Services 0.652 0.929 0.993 1.137 0.709 1.114 1.354 1.568 0.687 0.963 1.189 1.306
(0.148) (0.106) (0.079) (0.077) {0.596, 0.896} {0.983, 1.321} {1.196, 1.577} {1.357, 1.905} {0.577, 0.798} {0.843, 1.100} {1.052, 1.345} {1.136, 1.508}
@@
Manufacturing 0.969 1.092 0.962 1.062 0.545 1.577 1.959 2.231 0.592 1.398 1.836 2.039
(0.349) (0.370) (0.264) (0.294) {0.215, 1.015} {1.226, 2.101} {1.515, 2.549} {1.635, 3.268} {0.161, 0.996} {1.083, 1.853} {1.447, 2.366} {1.542, 2.903}
Utilities 0.277 0.581 0.751 0.987 0.518 1.538 1.632 2.137 0.545 1.540 1.743 2.222
(0.376) (0.302) (0.126) (0.217) {0.088, 1.046} {1.053, 4.050} {1.175, 2.760} {1.492, 4.366} {0.180, 0.929} {1.152, 2.289} {1.329, 2.498} {1.643, 3.481}
Construction 0.283 0.409 0.465 0.305 1.055 2.650 2.460 3.042 1.041 2.340 2.156 2.672
(0.640) (0.716) (0.304) (0.319) {–0.031, 2.979} {1.352, 5.528} {1.208, 4.698} {1.705, 5.782} {–0.063, 2.410} {0.985, 4.474} {1.089, 3.820} {1.571, 4.577}
Services 0.492 0.588 0.784 0.783 0.623 1.452 1.673 2.201 0.655 1.441 1.668 2.152
(0.348) (0.253) (0.114) (0.146) {0.303, 1.013} {1.174, 2.022} {1.392, 2.234} {1.719, 3.424} {0.349, 0.928} {1.176, 2.010} {1.403, 2.158} {1.716, 3.061}

### v9

@@ 3.4 Effects of technological change
This equation is a multi-sector, multi-factor general equilibrium version of the Krusell, Ohanian, Ríos-Rull and Violante (2000) model. It implies that the direction and magnitude of the effect of capital-embodied technological progress on the skill wage gap depend on the presence and degree of capital–skill complementarity. The skill wage gap can increase with capital-embodied technological progress if capital equipment is more complementary to skilled labor than to unskilled labor (i.e., $\epsilon_{\ell_{u}k_{e}}^{\mathcal{F}}>\epsilon_{\ell_{h}k_{e}}^{\mathcal{F}}$). The rate of the increase becomes greater as the difference between the two elasticities or the income share of capital equipment increases. The effect of sector-specific disembodied technological change on the skill wage gap can be divided into two parts. The first part in the square bracket takes a similar form to the effect of capital-embodied technological change and captures the effect through a change in capital equipment. However, this has the opposite sign to that of the first term because sector-specific disembodied technological progress reduces the need for the use of capital equipment. The second part is related to the direct effect denoted by $\psi^{\ell}$ in equation (18). The sign of the first part is negative in the presence of capital–skill complementarity, whereas that of the second part is likely to be positive. In this setting, sector-specific disembodied technological progress is unlikely to account for the widening of the skill wage gap.
@@ 4 Data
We have shown in Propositions 1 to 3 that the aggregate elasticities of substitution and the factor shares of income determine the direction and magnitude of changes in relative wages and income shares of different types of labor attributable to (i) changes in specific inputs, under which other inputs are held constant, and (ii) changes in technologies, under which all inputs are allowed to vary. This section begins by describing the data sources, sample, and variables used to calculate the factor shares of income and to estimate the aggregate elasticities of substitution. It then documents trends in the relative prices, relative quantities, and income shares of factors. The section ends by looking at the substitutability between ICCT capital and the four types of labor.
@@ 4.1 Data sources, sample, and variables
The data used in the analysis are mainly from the EU KLEMS database, which collects detailed and internationally comparable information on the prices and quantities of capital and labor in major OECD countries. Our analysis is based on the March 2008 version because it contains the longest time series covering the 1980s, during which there were dramatic changes in technology and inequality. For the years 1980 to 2005, we include all countries and years in the sample for which required data are available. Our sample comprises 11 OECD countries: Australia, Austria, the Czech Republic, Denmark, Finland, Germany, Italy, Japan, the Netherlands, the United Kingdom, and the United States. The economy of each country is divided into five sectors: primary, manufacturing, utilities, construction, and services. The sample comprises four non-primary sectors and includes 258 country-year observations for each sector. [^7] On average in our sample, the male share of employment is higher in the construction and utilities sectors than in any subsector of manufacturing or services.
Labor is divided into skilled and unskilled labor, each of which is further divided into male and female labor. Skilled labor consists of workers who have completed college, and unskilled labor consists of workers who do not enter or complete college. We calculate wages by gender and skill by dividing the total labor compensation by the total hours worked. When we calculate wages and hours worked, we adjust for changes in the age and education composition of the labor force over time. Appendix D.1 provides the details of the adjustment procedure.
Capital is divided into ICCT and non-ICCT capital. ICCT capital consists of computing equipment, communications equipment, and software. Non-ICCT capital consists of transport equipment, other machinery and equipment, non-residential structures and infrastructures, residential structures, and other assets. We calculate the rental price of capital (Jorgenson, 1963), based on either the internal or external rate of return. The former approach has the advantage of maintaining consistency between national income and production accounts, whereas the latter approach has the advantage of not requiring the assumption of competitive markets. We employ the latter approach to examine the robustness of our results. Appendix D.2 provides the details of the calculation procedure.
@@ 4.2 Factor prices, quantities, and shares
During the period between the years 1980 and 2005, there was a difference in the trends of the gender wage gap between skilled and unskilled workers in OECD countries (Figure 1a). The male–female wage gap declined among unskilled workers, but not among skilled workers. At the same time, there was a difference in the trends of the skill wage gap between male and female workers in OECD countries (Figure 1b). The skilled–unskilled wage gap increased among male workers after a slight drop in the early 1980s, while it did not increase among female workers except for the period between the years 1987 and 1995. Taken together, the rate of decrease in the male–female wage gap was greater among unskilled workers than among skilled workers, while the rate of increase in the skilled–unskilled wage gap was greater among male workers than among female workers. This finding is perhaps surprising but not inconsistent with the fact that the rate of decline in the gender wage gap was greater at the middle or bottom than at the top of the wage distribution in the United States from the 1980s to the 2000s (Blau and Kahn, 2017). The patterns of changes in gender and skill premia are more visible in the United States (Figures 2a and 2b). [^9] Since the mid-2000s, the widening of the skill wage gap has largely stabilized, while the narrowing of the gender wage gap has decelerated (Heathcote, Perri, Violante and Zhang, 2023). Across most countries in our sample, the rate of decline in the gender wage gap tends to be greater among unskilled workers than among skilled workers, while the rate of increase in the skill wage gap tends to be greater among male workers than among female workers. [^10] This pattern is observed in Australia, the Czech Republic, Denmark, Germany, Japan, and the United States. For the rest of the countries, there is no notable difference in the magnitude of changes in the gender wage gap between skilled and unskilled workers and in the skill wage gap between male and female workers.
@@
Turning to the relative quantities of labor, there was a decline in the quantity of male labor relative to female labor, and an increase in the quantity of skilled labor relative to unskilled labor. The rate of decline in the relative quantity of male labor was much greater for skilled labor than for unskilled labor (Figure 3a). At the same time, the rate of increase in the relative quantity of skilled labor was much greater for female labor than for male labor (Figure 3b). These observations suggest that the relative supply of labor alone could not explain the differences in the trends of the gender wage gap across skill groups and of the skill wage gap across gender groups.
@@
Figure 4: ICCT and non-ICCT capital
(a) Investment prices
@@
Notes: The investment prices of ICCT and non-ICCT capital are denoted by $p_{i}$ and $p_{o}$, respectively, and the output price is denoted by $p$. The quantities of ICCT and non-ICCT capital are denoted by $k_{i}$ and $k_{o}$, respectively. All the series are logarithmically transformed and normalized to zero in the year 1980.
The trends in the rental prices of capital differ significantly between ICCT and non-ICCT capital (Figure 4a). The rental price of ICCT capital fell dramatically, but that of non-ICCT capital remained almost unchanged. Meanwhile, the quantities of both ICCT and non-ICCT capital increased. However, the rate of increase in ICCT capital was far greater than that in non-ICCT capital (Figure 4b). These observations indicate substantial progress in ICCT during the period. [^11] Since the mid-2000s, the deepening of ICT capital has slowed markedly, while the decline in ICT investment prices has flattened considerably (Byrne and Corrado, 2017; Sichel, 2019; Goldin, Koutroumpis, Lafond and Winkler, 2024). Appendix D.3 shows that even if non-ICCT equipment is distinguished from non-ICCT structures, there is no significant difference in the trends of the prices and quantities between them.
The direction and magnitude of changes in labor shares vary by gender and skill (Figure 5). The income shares of male and female unskilled labor decreased, while the income shares of male and female skilled labor increased. The magnitude of the decline in the income share of male unskilled labor is substantially greater than that of the changes in the income shares of the other three types of labor. Taken together, the decline in the labor share is attributable to the fall in the income share of male unskilled labor. This finding is consistent with evidence of a decline in the income share of routine labor (Eden and Gaggl, 2018).
@@ 4.3 Factor substitution
Before we specify the sectoral production function, we conduct a preliminary analysis to acquire a general idea of substitutability between ICCT capital and the four types of labor. For labor of type $f$ in sector $n$, we estimate the following equation:
@@
where $\alpha_{nj}$ is a country-specific unobserved effect, $\nu_{njt}$ is the error term, and $j$ and $t$ are indices for countries and years. The parameter of interest is $\varepsilon_{f,n}$, which captures the percentage change in the capital–labor ratio associated with a one percent change in the relative price of labor to capital. The greater $\varepsilon_{f,n}$ is, the more substitutable labor of type $f$ is with ICCT capital in sector $n$.
We address two econometric issues: (i) unobserved heterogeneity and (ii) simultaneity. First, we control for country-specific unobserved effects by first differencing. In doing so, we take 5-year differences to deal with the measurement error problem. Then, we make use of two supply shifters as instrumental variables to identify equation (23) regarding the relative demand for factors. The first instrument is the lagged population size of each demographic group. The second instrument is the lagged price of ICCT investment relative to consumption, which measures investment-specific technology accelerating capital accumulation.
Table 1: Substitutability between ICCT capital and the four types of labor
OLS First-difference GMM1 First-difference GMM2
@@
Manufacturing 0.737 1.092 1.135 1.225 0.592 1.182 1.537 1.613 0.562 0.986 1.305 1.339
(0.136) (0.144) (0.179) (0.154) {0.378, 0.829} {1.041, 1.364} {1.298, 1.823} {1.392, 1.893} {0.373, 0.751} {0.832, 1.160} {1.050, 1.607} {1.113, 1.604}
Utilities 0.369 0.869 0.941 1.207 0.472 1.088 1.283 1.513 0.482 0.975 1.156 1.319
(0.234) (0.159) (0.058) (0.143) {0.321, 0.689} {0.841, 1.513} {1.044, 1.683} {1.235, 2.009} {0.347, 0.647} {0.696, 1.339} {0.886, 1.504} {1.029, 1.692}
Construction 0.728 1.016 1.015 1.004 1.107 1.828 1.819 2.027 0.923 1.373 1.456 1.583
(0.338) (0.445) (0.201) (0.254) {0.413, 2.047} {1.109, 2.865} {1.160, 2.753} {1.368, 3.024} {0.454, 1.558} {0.874, 2.053} {0.970, 2.123} {1.092, 2.277}
Services 0.652 0.929 0.993 1.137 0.713 1.109 1.350 1.554 0.687 0.963 1.189 1.306
(0.148) (0.106) (0.079) (0.077) {0.592, 0.858} {0.978, 1.302} {1.193, 1.573} {1.344, 1.886} {0.577, 0.798} {0.843, 1.100} {1.052, 1.345} {1.136, 1.508}
@@
Manufacturing 0.969 1.092 0.962 1.062 0.599 1.564 1.967 2.232 0.592 1.398 1.836 2.039
(0.349) (0.370) (0.264) (0.294) {0.199, 1.009} {1.200, 2.109} {1.531, 2.545} {1.650, 3.247} {0.161, 0.996} {1.083, 1.853} {1.447, 2.366} {1.542, 2.903}
Utilities 0.277 0.581 0.751 0.987 0.490 1.397 1.586 2.047 0.545 1.540 1.743 2.222
(0.376) (0.302) (0.126) (0.217) {0.090, 1.000} {0.899, 3.007} {1.110, 2.709} {1.401, 4.071} {0.180, 0.929} {1.152, 2.289} {1.329, 2.498} {1.643, 3.481}
Construction 0.283 0.409 0.465 0.305 1.137 2.712 2.474 3.046 1.041 2.340 2.156 2.672
(0.640) (0.716) (0.304) (0.319) {0.023, 2.622} {1.292, 5.372} {1.251, 4.577} {1.737, 5.588} {–0.063, 2.410} {0.985, 4.474} {1.089, 3.820} {1.571, 4.577}
Services 0.492 0.588 0.784 0.783 0.662 1.450 1.668 2.170 0.655 1.441 1.668 2.152
(0.348) (0.253) (0.114) (0.146) {0.339, 0.950} {1.172, 2.079} {1.388, 2.254} {1.695, 3.397} {0.349, 0.928} {1.176, 2.010} {1.403, 2.158} {1.716, 3.061}